# No.2 Linggong Road, Dalian, Liaoning, 116024, P.R.China

### v1

@@ 2.3 Learnable cut flow
The model takes the original data as inputs. Then data is normalized with the mean and standard deviation of the training data. Then data is split along the observable axis to feed into each learnable cut. The output of each learnable cut is collected at the end. When the LCF model performs inference on new data, a minimum importance ratio is applied, set to 0.05 by default. This means it automatically ignores cuts with observables below 5% of the average, where the average is defined as 1 divided by the number of features.
@@
$L_{ij}=\frac{1}{2}\left(L^{\text{lower}}_{ij}+L^{\text{upper}}_{ij}\right)\cdot\text{Mask}_{i,j-1}$ (35)
@@
$\text{Mask}_{i,j}=\begin{cases}1&\text{if }j=0\\ \Theta(\hat{y}^{\text{lower}}_{i,j-1}-t)\cdot\Theta(\hat{y}^{\text{upper}}_{i,j-1}-t)&\text{if }j>0\end{cases}$ (36)
@@ 3.1 Datasets
This study employs two types of datasets: a series of synthetic mock datasets to demonstrate the effectiveness of the proposed neural network under controlled and simplified conditions, and a real dataset from 1603.09349 to benchmark its performance against baseline models.
@@
$x_{1}$ $N(-2,4)$ $N(2,4)$ Left case
$x_{2}$ $N(2,4)$ $N(-2,4)$ Right case
$x_{3}$ $N(0,4)$ $N(5,4)+N(-5,4)$ Middle case
$x_{4}$ $N(5,4)+N(-5,4)$ $N(0,4)$ Edge case
$x_{5}$ $N(-1,4)$ $N(1,4)$ Weakly separated
$x_{6}$ $N(-3,4)$ $N(3,4)$ Strongly separated
$x_{7}$ $N(-5,4)$ $N(-5,4)$ Redundant
@@
Figure 2: Mock feature distributions.
@@
  • Mock4 combines $x_{1}$ with two highly correlated features ($x_{9}$, $x_{10}$) to simulate a common challenge in phenomenology data analysis.
@@
  • Mock6 randomly reorders the features of Mock5 to evaluate the model’s sensitivity to feature order.
Figure 3: Feature correlations of mock datasets.
The real dataset was created using standard simulation tools in high energy physics. The signal, diboson production $pp\rightarrow W^{+}W^{-}\rightarrow qqqq$, produces two fat jets exhibiting a 2-prong substructure. In contrast, the background, QCD dijet production $pp\rightarrow qq,qg,gg$ predominantly yields 1-prong jets.
Events were generated using MadGraph 5 v2.2.3 for hard scattering, Pythia v6.426 for showering and hadronization, and Delphes v3.2.0 for detector response. Jets were reconstructed with the anti-$k_{t}$ algorithm and parameter $R=1.2$ using FastJet v3.1.2. Their constituents were then reclustered with the $k_{t}$ algorithm and parameter $R=0.2$ to form subjets. Subjets with $p_{T}^{\text{sub}}<p_{T}^{\text{jet}}\times 3\%$ were discarded and the remaining subjets were used to construct the trimmed jets. To ensure the collinear $W$ bosons decays, only jets with $p_{T}^{\text{trim}}\in[300,400]$ GeV were retained.
@@
Figure 4: Distributions of jet substructure variables. Top row: jet mass (left) and $\tau_{21}$ with $\beta=1$ (right). Middle row: $C_{2}$ with $\beta=1$ (left) and $\beta=2$ (right). Bottom row: $D_{2}$ with $\beta=1$ (left) and $\beta=2$ (right).
@@
Both datasets contains 200,000 samples evenly divided between signal and background. The data are split into a training set (50%) and a test set (50%). To maintain computational efficiency, samples of the real dataset are randomly drawn from the original non-pile up test set. Furthermore, the original observable distributions show a significant presence of low values for all observables except $M_{\text{jet}}$. To mitigate outliers and ensure stable training, only samples within the 5th to 95th percentile range of all observables are retained. Both dataset types define a binary classification task, allowing the learnable cut flow models to be well trained and provide insights into the neural network’s decision-making process.
3.2 Baseline models
@@
  • The BDT, built with Scikit-Learn, consists of 100 estimators with a maximum depth of 3. It is trained directly on raw input data without preprocessing as it is insensitive to feature scaling. The output represents the probability of the signal event.
  • The MLP, built with Keras, includes a standard normalization layer for input preprocessing and five hidden layers with 16, 32, 64, 32, and 16 units respectively. All hidden layers use the rectified linear unit (ReLU) activation function. The output layer applies a sigmoid function to produce the signal probability.
All models involved (BDT, MLP, LCF in two strategies) are trained using the same configuration: binary cross entropy as the loss function, Adam as the optimizer with a learning rate 0.001, a batch size of 512, and 200 training epochs. Early stopping is not applied as training of LCF exhibits plateau phases in loss. The centers for LCF model are set as -2, 2, 0, 0, -1, -3, -5, 5, -1.8, -1.4 for each mock feature, and 80, 0.15, 0.025, 2, 2, 0.3 for each observable in the diboson dataset.
@@ 4.1 Learned cuts
In the parallel strategy, the model analyzes each feature’s distribution independently to optimize its cut positions. As shown in figure 6, the resulting cuts effectively isolate signal-rich regions: $x_{1}$ captures the left tail of the signal, $x_{2}$ the right tail, $x_{3}$ the central peak, and $x_{4}$ the outer tails. These positions align well with the boundaries designed for Mock1. By contrast, the sequential strategy applies cuts iteratively, modifying the distribution after each cut to account for inter-feature correlations. Figure 7 illustrates how the cut positions adapt to these updates, with subsequent cuts refining the remaining signal region to improve signal-background separation as the feature space is progressively filtered by earlier cuts.
The importance scores shown in figure 9 and figure 9 reflect how the LCF prioritizes each feature. In the parallel strategy, scores are relatively balanced: features with single-sided cuts ($x_{1}$ and $x_{2}$) have scores 0.179 and 0.185, respectively, compared to features with double-sided cuts ($x_{3}$ and $x_{4}$) at 0.135 and 0.321. By contrast, the importance shifts in the sequential strategy as prior cuts filter out a fraction of background events: $x_{1}$ increases to 0.224 while $x_{2}$ decreases to 0.164; similarly, $x_{3}$ increases to 0.351 and $x_{4}$ decreases to 0.260.
@@
Figure 6: Learned cuts from the parallel LCF on the Mock1 dataset.
Figure 7: Learned cuts from the sequential LCF on the Mock1 dataset.
Figure 8: Learned importance from the parallel LCF on the Mock1 dataset.
Figure 9: Learned importance from the sequential LCF on the Mock1 dataset.
@@ 4.2 Learned importance
To further evaluate the LCF model’s ability to prioritize features, we analyze the importance scores learned on the Mock2 dataset, which includes three features ($x_{1}$, $x_{5}$, $x_{6}$) with regular, weak, and strong signal-background separation. The signal and background distributions, along with the learned cuts, are shown in figure 10 (parallel LCF) and figure 11 (sequential LCF). The corresponding importance scores are shown in figure 13 and figure 13.
In the parallel strategy, the LCF model assigns importance scores that closely match each feature’s intended role: compared to $x_{1}$’s regular signal-background separation, $x_{5}$ decrease it while $x_{6}$ increase it. In the sequential strategy, however, cuts modify the features space step by step, which reduces $x_{5}$’s effective separation and thus lowers its importance. Meanwhile, $x_{6}$ maintains a clear boundary between signal and background, resulting in a score similar to $x_{1}$.
These results demonstrate that the learnable importance mechanism captures each features’ discriminative power and adapts dynamically under both the parallel and sequential strategies.
Figure 10: Learned cuts from the parallel LCF on the Mock2 dataset.
Figure 11: Learned cuts from the sequential LCF on the Mock2 dataset.
Figure 12: Learned importance from the parallel LCF on the Mock2 dataset.
Figure 13: Learned importance from the sequential LCF on the Mock2 dataset.
@@ 4.3 Robustness to redundancy
To assess how robust the LCF models are when presented with redundant features, we train them on the Mock3 dataset. In this setup, $x_{1}$ provides a clear signal-background separation, whereas $x_{7}$ and $x_{8}$ offer little to no additional discriminative power. As shown in figure 14, the learned cuts for $x_{7}$ and $x_{8}$ appear almost randomly placed, reflecting their limited ability to distinguish signal and background. The corresponding importance scores in figure 17 confirm that the model recognizes these two features contribute minimally to separation.
In the sequential strategy, the first cut leads to a nearly complete overlap of signal and background for $x_{7}$ and $x_{8}$ as shown in figure 15. This, in turn, results in almost zero importance for these features (see figure 17).
@@
Figure 14: Learned cuts from the parallel LCF on the Mock3 dataset.
Figure 15: Learned cuts from the sequential LCF on the Mock3 dataset.
Figure 16: Learned importance from the parallel LCF on the Mock3 dataset.
Figure 17: Learned importance from the sequential LCF on the Mock3 dataset.
@@ 4.4 Robustness to high correlation
High correlation between observables implies that a cut on one feature effectively constrains the others, posing a challenge for cut-based methods. To evaluate the LCF models’ robustness in this scenario, we train them on the Mock4 dataset which includes $x_{1}$ alongside highly correlated features $x_{9}$ and $x_{10}$. The resulting learned cuts are shown in figure 18 (parallel LCF) and figure 19 (sequential LCF), with corresponding importance scores shown in figure 21 and figure 21.
Figure 18: Learned cuts from the parallel LCF on the Mock4 dataset.
Figure 19: Learned cuts from the sequential LCF on the Mock4 dataset.
Figure 20: Learned importance from the parallel LCF on the Mock4 dataset.
Figure 21: Learned importance from the sequential LCF on the Mock4 dataset.
In the parallel strategy, the LCF model treats each feature independently, so the high correlation doesn’t significantly affect the learned cuts or importance scores. These scores remain consistent with the dataset’s design. By contrast, since the first cut in the sequential strategy reshapes the subsequent feature distributions, $x_{9}$ and $x_{10}$ end up with lower importance.
Both strategies successfully identify the proper cut boundaries under high correlation, demonstrating that the LCF models remain robust in this setting.

### v3

@@ 2.3 Learnable cut flow
The model takes the original data as inputs. Then data are normalized with the mean and standard deviation of the training data. Then data are split along the observable axis to feed into each learnable cut. The output of each learnable cut is collected at the end. When the LCF model performs inference on new data, a minimum importance ratio is applied, set to 0.05 by default. This means it automatically ignores cuts with observables below 5% of the average (defined as the baseline in the context of the learned importance), where the average is defined as 1 divided by the number of features.
@@
$L_{ij}=\frac{1}{2}\left(L^{\text{lower}}_{ij}+L^{\text{upper}}_{ij}\right)\cdot\text{Mask}_{i,j-1},$ (35)
@@
$\text{Mask}_{i,j}=\begin{cases}1&\text{if }j=0\\ \Theta(\hat{y}^{\text{lower}}_{i,j-1}-t)\cdot\Theta(\hat{y}^{\text{upper}}_{i,j-1}-t)&\text{if }j>0\end{cases}.$ (36)
@@ 3.1 Datasets
This study employs two types of datasets: a series of synthetic mock datasets to demonstrate the effectiveness of the proposed neural network under controlled and simplified conditions, and a real dataset produced in the context of the studies presented in ref. 1603.09349 to benchmark its performance against baseline models.
@@
$x_{1}$ N(2,4)N(-2,4) $N(2,4)$ Left case
$x_{2}$ $N(2,4)$ N(2,4)N(-2,4) Right case
$x_{3}$ $N(0,4)$ N(5,4)+N(5,4)N(5,4)+N(-5,4) Middle case
$x_{4}$ N(5,4)+N(5,4)N(5,4)+N(-5,4) $N(0,4)$ Edge case
$x_{5}$ N(1,4)N(-1,4) $N(1,4)$ Weakly separated
$x_{6}$ N(3,4)N(-3,4) $N(3,4)$ Strongly separated
$x_{7}$ N(5,4)N(-5,4) N(5,4)N(-5,4) Redundant
@@
Figure 2: Mock feature distributions from $x_{1}$ to $x_{10}$. $x_{1}$, $x_{2}$, $x_{3}$, and $x_{4}$ represent four basic cases of signal relative location. $x_{5}$ and $x_{6}$ have weaker and stronger separation compared with $x_{1}$. $x_{7}$ and $x_{8}$ are two redundant features. $x_{9}$ and $x_{10}$ are two features highly-correlated with $x_{1}$. The center value divides a distribution into two parts to train two sets of trainable weights to form one learned cut.
@@
  • Mock4 combines $x_{1}$ with two highly correlated features ($x_{9}$, $x_{10}$) to simulate a common challenge in data analysis, where correlated observables often complicate signal-background separation.
@@
  • Mock6 randomly reorders the features of Mock5 to evaluate the model’s sensitivity to feature order. The order of features is $x_{1},x_{2},x_{7},x_{3},x_{5},x_{4},x_{9}$ since the random seed is fixed.
(a) Mock1
(b) Mock2
(c) Mock3
(d) Mock4
(e) Mock5
(f) Mock6
Figure 3: Feature correlations of mock datasets. Mock4 dataset contains two features $x_{9}$ and $x_{10}$ highly-correlated to $x_{1}$.
The real dataset was created using standard simulation tools in high energy physics. The signal, diboson production $pp\rightarrow W^{+}W^{-}\rightarrow qqqq$, produces two fat jets exhibiting a 2-prong substructure. In contrast, the background, QCD dijet production $pp\rightarrow qq,qg,gg,$ predominantly yields 1-prong jets.
Events were generated using MadGraph 5 v2.2.3 for hard scattering, Pythia v6.426 for showering and hadronization, and Delphes v3.2.0 for detector response. Jets were reconstructed with the anti-$k_{t}$ algorithm and parameter $R=1.2$ using FastJet v3.1.2. Their constituents were then reclustered with the $k_{t}$ algorithm and parameter $R=0.2$ to form subjets. Subjets with $p_{T}^{\text{sub}}<p_{T}^{\text{jet}}\times 3\%$ were discarded and the remaining subjets were used to construct the trimmed jets 0912.1342 . To ensure the collinear $W$ bosons decays, only jets with $p_{T}^{\text{trim}}\in[300,400]$ GeV were retained.
@@
Figure 4: Distributions of jet substructure variables in the real diboson vs. QCD dataset. The center value divides a distribution into two parts to train two sets of trainable weights to form one learned cut.
@@
Each dataset contains 200,000 events evenly divided into signal and background. The data are split into a training set (50%) and a test set (50%). The real dataset, sourced from 1603.09349 , is originally split into training and test sets, each further divided into pile-up and non-pileup subsets. To ensure computational efficiency, we didn’t use the full sets; instead, our events are randomly drawn from the original non-pileup test set to reduce computational load while maintaining representativity. Furthermore, the original observable distributions show a significant presence of low values for all observables except $M_{\text{jet}}$. This is a physical effect that arises from the one-pronged nature of QCD background jets. Unlike signal jets from diboson decays, which exhibit multi-pronged substructure, QCD jets typically originate from single quarks or gluons and retain this characteristic even when clustered using algorithms with relatively large R parameters. For further details on the distributions of specifically selected dijet QCD events, see refs. 1011.2268 ; 1305.0007 . To mitigate outliers and ensure stable training, only events within the 5th to 95th percentile range of all observables are retained. Both dataset types define a binary classification task, allowing the learnable cut flow models to be well trained and provide insights into the neural network’s decision-making process.
3.2 Baseline models and training
@@
  • The BDT, built with Scikit-Learn scikit-learn , consists of 100 estimators with a maximum depth of 3. It is trained directly on raw input data without preprocessing as it is insensitive to feature scaling. The output represents the probability of the signal event.
  • The MLP, built with Keras keras , includes a standard normalization layer for input preprocessing and five hidden layers with 16, 32, 64, 32, and 16 units respectively. All hidden layers use the rectified linear unit (ReLU) activation function. The output layer applies a sigmoid function to produce the signal probability.
All models involved (BDT, MLP, LCF in two strategies) are trained using the same configuration: binary cross entropy as the loss function, Adam as the optimizer with a learning rate 0.001, a batch size of 512, and 200 training epochs. Early stopping is not applied as training of LCF exhibits plateau phases in loss. The centers for the LCF model are set as -2, 2, 0, 0, -1, -3, -5, 5, -1.8, -1.4 for each mock feature as shown in figure 2 and 80, 0.15, 0.025, 2, 2, 0.3 for each observable in the diboson dataset as shown in figure 4. Each center is chosen to divide the respective observable’s distribution into two regions, enabling the LCF to learn boundaries on one side (for left or right cases) or both sides (for middle or edge cases). The choice of center values is not strict. For single-sided cuts, the center can be placed anywhere except at the expected boundary. For double-sided cuts, it is ideally positioned near the middle of the distribution to enable learning of boundaries on both sides. If misplaced—e.g., at the distribution’s leftmost edge—the double-sided cuts will degenerate into single-sided cuts, mismatching the distribution and reducing performance. Models are trained continuously ten times to demonstrate robustness and stability.
3.3 Performance evaluation
To comprehensively assess the performance of the LCF model and compare it with baseline models, we employ the following metrics, evaluated on the test set, that capture different aspects of classification performance. For each metric, higher values are preferable unless otherwise noted:
  • True Positive (TP): The number of simulated signal events correctly identified (i.e., retained by the cuts).
  • False Positive (FP): The number of simulated background events incorrectly identified as signal (i.e., passing the cuts).
  • Accuracy: The fraction of events (both signal and background) that are correctly classified. It’s calculated as (TP + TN) / total events, where TN is true negatives (the number of background events correctly identified).
  • Precision: The ratio of correctly identified signal events to all events predicted as signal, calculated as TP / (TP + FP).
  • Significance: A measure of sensitivity to a rare signal against a dominant background in high-energy physics. It’s computed as $S/\sqrt{B}$, where $S=\epsilon_{s}\times\sigma_{s}\times\mathcal{L}$ is the expected number of signal events and $B=\epsilon_{b}\times\sigma_{b}\times\mathcal{L}$ is the expected number of background events after all learned cuts. Here, $\epsilon_{s}=$TP / total test signal events is the signal efficiency, $\epsilon_{b}=$FP / total test background events is the background efficiency. The integrated luminosity $\mathcal{L}$ is assumed to be 3000 $\text{fb}^{-1}$. For the mock datasets, we assume: $\sigma_{s}=1$ pb (rare signal) and $\sigma_{b}=10^{6}$ pb (abundant background). For the real diboson dataset, we estimate cross sections based on the original processes: approximately 0.7644 pb for the diboson signal and $1.806\times 10^{5}$ pb for the QCD background. Since ref. 1603.09349 does not provide cross sections after their full processing pipeline, we derived cross sections independently using only anti-kt jet clustering ($R=1.2$). These nominal cross sections are used for basic significance comparisons and differ from those in the reference due to the omitted trimming and selection steps.
@@ 4.1 Learned cuts
Figure 6: Learned cuts from the parallel LCF on the Mock1 dataset. The error bars on the learned cut boundaries are negligible compared to the bin width and are therefore not shown.
Figure 7: Learned cuts from the sequential LCF on the Mock1 dataset. The error bars on the learned cut boundaries are negligible compared to the bin width and are therefore not shown.
@@
In the parallel strategy, the model analyzes each feature’s distribution independently to optimize its cut positions. As shown in figure 6, the resulting cuts effectively isolate signal-rich regions: $x_{1}$ captures the left tail of the signal, $x_{2}$ the right tail, $x_{3}$ the central peak, and $x_{4}$ the outer tails. These positions align well with the boundaries designed for Mock1. In contrast, the sequential strategy applies cuts iteratively, modifying the distribution after each cut to account for inter-feature correlations. Figure 7 illustrates how the cut positions adapt to these updates, with subsequent cuts refining the remaining signal region to improve signal-background separation as the feature space is progressively filtered by earlier cuts.
Figure 8: Learned importance from parallel (left) and sequential (right) LCFs on the Mock1 dataset. The baseline indicates the minimum importance (default: 5% of average importance = $1/F\times 0.05$), below which the features are ignored during inference.
The importance scores shown in figure 8 reflect how the LCF prioritizes each feature. The baseline shown in the same figure represents the minimum threshold (5% of average importance) below which features are automatically excluded during inference. In the parallel strategy, scores are relatively balanced: features with single-sided cuts ($x_{1}$ and $x_{2}$) have scores 0.1842 and 0.1814, respectively, compared to features with double-sided cuts ($x_{3}$ and $x_{4}$) at 0.3166 and 0.3179. In contrast, the importance shifts in the sequential strategy as prior cuts filter out a fraction of background events: $x_{1}$ increases to 0.2198 while $x_{2}$ decreases to 0.1685; similarly, $x_{3}$ increases to 0.3474 and $x_{4}$ decreases to 0.2642.
Model TP FP Accuracy Precision Significance
LCF (Parallel) $25734\pm 66$ $34\pm 0$ $75.8\pm 0.1\%$ $99.9\pm 0.0\%$ $34.3\pm 0.1$
LCF (Sequential) $39759\pm 23$ $1577\pm 19$ $88.3\pm 0.0\%$ $96.2\pm 0.0\%$ $7.8\pm 0.0$
Table 2: Performance metrics of LCF models on the Mock1 dataset.
In table 2, we observe that the sequential strategy captures more true signal events and achieves higher accuracy because it examines the distributions more carefully by considering the relationships among features. However, this strategy also retains more background events, resulting in lower precision and significance. The parallel LCF significantly reduces the number of background events at the cost of modest signal loss, achieving higher precision and remarkably high significance.
@@ 4.2 Learned importance
Figure 9: Learned cuts from the parallel LCF on the Mock2 dataset. The error bars on the learned cut boundaries are negligible compared to the bin width and are therefore not shown.
Figure 10: Learned cuts from the sequential LCF on the Mock2 dataset. The error bars on the learned cut boundaries are negligible compared to the bin width and are therefore not shown.
Figure 11: Learned importance from parallel (left) and sequential (right) LCFs on the Mock2 dataset. The baseline indicates the minimum importance (default: 5% of average importance = $1/F\times 0.05$), below which the features are ignored during inference.
To further evaluate the LCF model’s ability to prioritize features, we analyze the importance scores learned on the Mock2 dataset, which includes three features ($x_{1}$, $x_{5}$, $x_{6}$) with regular, weak, and strong signal-background separation. The signal and background distributions, along with the learned cuts, are shown in figure 9 (parallel LCF) and figure 10 (sequential LCF). The corresponding importance scores are shown in figure 11.
In the parallel strategy, the LCF model assigns importance scores that closely match each feature’s intended role: compared to $x_{1}$’s regular signal-background separation, $x_{5}$ decreases it while $x_{6}$ increases it. The sequential strategy doesn’t change much the rank of the features. The score of the earlier cut on $x_{1}$ is raised a lot and later cuts’ scores are lowered a little bit, which matches what we find in section 4.1.
Model TP FP Accuracy Precision Significance
LCF (Parallel) $27117\pm 58$ $155\pm 1$ $77.1\pm 0.1\%$ $99.4\pm 0.0\%$ $17.0\pm 0.0$
LCF (Sequential) $40553\pm 24$ $1275\pm 7$ $89.4\pm 0.0\%$ $97.0\pm 0.0\%$ $8.8\pm 0.0$
Table 3: Performance metrics of LCF models on the Mock2 dataset.
Table 3 shows the same pattern observed in section 4.1. These results demonstrate that the learnable importance mechanism captures each features’ discriminative power and adapts dynamically under both the parallel and sequential strategies.
@@ 4.3 Robustness to redundancy
Figure 12: Learned cuts from the parallel LCF on the Mock3 dataset. The error bars are not shown because $x_{1}$’s error is negligible compared to the bin width while $x_{7}$’s and $x_{8}$’s errors have no meaning when the importance is close to zero.
Figure 13: Learned cuts from the sequential LCF on the Mock3 dataset. The error bars are not shown because $x_{1}$’s error is negligible compared to the bin width while $x_{7}$’s and $x_{8}$’s errors have no meaning when the importance is close to zero.
Figure 14: Learned importance from parallel (left) and sequential (right) LCFs on the Mock3 dataset. The baseline indicates the minimum importance (default: 5% of average importance = $1/F\times 0.05$), below which the features are ignored during inference.
To assess how robust the LCF models are when presented with redundant features, we train them on the Mock3 dataset. In this setup, $x_{1}$ provides a clear signal-background separation, whereas $x_{7}$ and $x_{8}$ offer little to no additional discriminative power. As shown in figure 12 and figure 13, the learned cuts on $x_{7}$ and $x_{8}$ appear almost randomly placed, reflecting their limited ability to distinguish signal and background. The corresponding importance scores in figure 14 confirm that both parallel and sequential strategies identify these two features contribute minimally to separation.
Model TP FP Accuracy Precision Significance
LCF (Parallel) $38673\pm 8392$ $7443\pm 1624$ $81.4\pm 6.8\%$ $83.9\pm 0.1\%$ $3.5\pm 0.5$
LCF (Sequential) $41924\pm 21$ $8060\pm 21$ $84.0\pm 0.0\%$ $83.9\pm 0.0\%$ $3.6\pm 0.0$
Table 4: Performance metrics of LCF models on the Mock3 dataset.
Table 4 shows that both strategies achieve almost the same performance. Across ten training runs, $x_{7}$’s cuts in the parallel strategy show inconsistent importance scores relative to the baseline—while eight runs yield scores below the baseline, two runs obtain scores above it (0.0183 and 0.0187). When $x_{7}$’s importance exceeds the baseline, the cut on it is applied during inference; when below, it is ignored entirely. Such treatment leads to higher metric errors in the parallel strategy then those in the sequential strategy: accuracy error of $6.8\%$ vs $0.0\%$ and significance error of $0.5$ vs $0.0$. Despite this variability, the average importance of $x_{7}$ is similar between both strategies, leading to the final similarity in significance obtained with the two strategies.
@@ 4.4 Robustness to high correlation
Figure 15: Learned cuts from the parallel LCF on the Mock4 dataset. The error bars on the learned cut boundaries are negligible compared to the bin width and are therefore not shown.
Figure 16: Learned cuts from the sequential LCF on the Mock4 dataset. The error bars on the learned cut boundaries are negligible compared to the bin width and are therefore not shown.
Figure 17: Learned importance from parallel (left) and sequential (right) LCFs on the Mock4 dataset. The baseline indicates the minimum importance (default: 5% of average importance = $1/F\times 0.05$), below which the features are ignored during inference.
High correlation between observables implies that a cut on one feature effectively constrains the others, posing a challenge for cut-based methods. To evaluate the LCF models’ robustness in this scenario, we train them on the Mock4 dataset which includes $x_{1}$ alongside highly correlated features $x_{9}$ and $x_{10}$. The resulting learned cuts are shown in figure 15 (parallel LCF) and figure 16 (sequential LCF), with corresponding importance scores shown in figure 17.
In the parallel strategy, the LCF model treats each feature independently, so the high correlation doesn’t significantly affect the learned cuts or importance scores. These scores remain consistent with the dataset’s design. In contrast, since the first cut in the sequential strategy reshapes the subsequent feature distributions, $x_{9}$ and $x_{10}$ end up with lower importance. Although the learned cuts of the sequential strategy almost cover the full range of the features, the importance scores are still not zero. It is because the cuts on $x_{9}$ and $x_{10}$ generate reasonable losses compared to $x_{7}$ and $x_{8}$ in the Mock3 dataset.
Model TP FP Accuracy Precision Significance
LCF (Parallel) $35010\pm 19$ $4710\pm 9$ $0.4\pm 0.0\% 88$$8.1\pm 0.0\%$ $4.0\pm 0.0$
LCF (Sequential) $41903\pm 18$ $8044\pm 16$ $84.0\pm 0.0\%$ $83.9\pm 0.0\%$ $3.6\pm 0.0$
Table 5: Performance metrics of LCF models on the Mock4 dataset.
Combined with the results in sections 4.1, 4.2, and 4.3, we observe that the parallel strategy typically achieves higher precision and significance, while the sequential strategy achieves higher accuracy. This difference most likely arises from a fundamental consequence of the parallel strategy’s independent feature processing. For features $x_{1}$, $x_{2}$, $x_{3}$, $x_{4}$, $x_{5}$, $x_{6}$, $x_{9}$, and $x_{10}$, portions of the background have already been filtered out by other cuts operating in parallel. This causes the cross point of the signal and background distributions to shift outward relative to the signal distribution. Consequently, the cuts learned from the original distributions happen to be more stringent and more conservative than the ones learned from the distributions after applying cuts on previous variables. This makes the parallel strategy achieve better performance on the Mock1, Mock2, and Mock4 datasets where such cuts remove a large portion of the background at a fair cost of signal efficiency. However, in the Mock3 dataset, random cuts learned from $x_{7}$ and $x_{8}$ make the reduction of signal and background an almost even trade-off, which has a negative impact on the rare signal events that are far less frequent than the background events. So the sequential strategy achieves better performance on the Mock3 dataset. Similarly, in other datasets where the parallel strategy’s tighter cut boundaries may incur signal efficiency costs that outweigh the background reduction benefits, the sequential strategy would be expected to achieve better significance.
Both strategies in the Mock4 dataset successfully identify the proper cut boundaries under high correlation, demonstrating that the LCF models remain robust in this setting.